# Beyond boundaries: Gary Lorden’s groundbreaking contributions to sequential analysis

### v1

@@ 2.6 Near Uniform Optimality of the GLR SPRT for Composite Hypotheses
and where $\zeta(\theta,\theta_{i})$, $i=0,1$, are defined in (51)–(52). These approximations are important for frequentist problems, which are of the main interest in most applications. Although there are no strict upper bounds on the error probabilities so that there is no prescription on how to embed the GSLRT into class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$, the asymptotic approximations (56) allow us to select thresholds $a_{i}$ in the stopping times $\widehat{T}_{i}$ so that $\hat{\alpha}_{i}(\theta_{i})\approx\alpha_{i}$, $i=0,1$, at least for sufficiently small $\alpha_{i}$. Note that in this latter case, $a$ in (53) should be replaced with $a_{i}$, the roots of the transcendental equations
@@
where the $o(1)$ term can be shown to be of order $O(\log|\log\alpha_{\max}|/|\log\alpha_{\max}|)$. We should note that this result is correct not only in the asymptotically symmetric case where $\log\alpha_{0}\sim\log\alpha_{1}$ and $a_{0}\sim a_{1}$ as $\alpha_{\max}\to 0$, but also in the asymmetric case where $a_{0}$ and $a_{1}$ go to infinity with different rates as long as $a_{1}e^{-a_{0}}\to 0$.
Note that the Schwarz–Lorden asymptotic theory assumes a fixed indifference zone that does not allow for local alternatives, i.e., $\theta_{1}$ cannot approach $\theta_{0}$ as $c\to 0$. In other words, this theory is limited to the case where the width of the indifference zone $\theta_{1}-\theta_{0}$ is considerably larger than $c^{1/2}$.
@@ 2.7 Optimal Multistage Testing
Beginning with the simple vs. simple testing setup of Section 2.2 and adopting the notation there, some of Lorden’s main ideas can be seen by first considering the symmetric case where the error probabilities $\alpha_{0},\alpha_{1}\to 0$ in such a way that
@@
$\frac{1/n}{(1-t/n)^{2}}=\frac{n}{\delta_{t}^{2}}=\frac{t+\delta_{t}}{\delta_{t}^{2}}=o(1)$
as $t\to\infty$. A similar argument shows that the other probability in (58) approaches 00 as well.
In this symmetric situation an optimal 2-stage competitor to the SPRT can be described in terms of $n(t)$, which is the size of the first stage with $t$ taken to be the larger of the two sides of (57). Note that, for either $i=0$ or $1$, we have $tI_{i}\sim\log\alpha_{1-i}^{-1}$ so that $t$ is asymptotically the same as the expected stopping time of the SPRT (under either hypothesis) and the first stage $n(t)$ is of the same order but slightly larger. The procedure stops after the first stage if
@@
Lorden shows that no 2-stage test can be asymptotically optimal; this itself is a nontrivial result which we discuss in the next section. But for this case Lorden gives a 3-stage procedure that is a slight modification of the one above. Letting $t_{1}$ and $t_{2}$ be the left- and right-hand sides of (57), respectively, the first stage of Lorden’s procedure is of size $\min\{n(t_{1}),n(t_{2})\}$, and the second stage (if needed) brings the total sample size to $\max\{n(t_{1}),n(t_{2})\}$, both using the stopping rule (60) and corresponding decision rule. If not stopped after the second stage, a third stage is employed that brings the total sample size to that of the fixed-sample size with error probabilities $\alpha_{i}$, which is $\leqslant C\max\{n(t_{1}),n(t_{2})\}$as above, and uses that terminal decision rule. Since
$n(t_{i+1})\sim t_{i+1}\sim\frac{\log\alpha_{1-i}^{-1}}{I_{i}}\quad\mbox{for both}\quad i=0,1$
and (61) and (62) hold for $n(t_{1})$ and $n(t_{2})$, respectively, the expected sample size of this 3-stage procedure is asymptotically equal to the corresponding side of (57), and is thus minimized under both the null and alternative.
@@ 2.7.2 The Necessity of 3 Stages
For testing $f_{0}$ vs. $f_{1}$ in the setup of Section 2.2, let $N$ denote the sample size of a $k$-stage ($k\geqslant 2$) test with error probabilities $\alpha_{0}$ and $\alpha_{1}$, and let $M$ be the total sample size of this test after $k-1$ stages. If $N$ is asymptotically optimal as $\alpha_{0},\alpha_{1}\to 0$ and $g$ is a density distinct from $f_{0}$ such that
@@
[…]
${\mathsf{E}}_{g}[M]$ $\sim\frac{\log\alpha_{0}^{-1}}{I(g,f_{0})}\sim{\mathsf{E}}_{g}[N]$
as $\alpha_{0}\to 0$.
Lorden’s proof of this theorem is technical and requires detailed upper bounds on the conditional error probabilities after the $(k-1)$st stage; that is, the probabilities of test error given the first $M$ observations. Roughly speaking, showing that these conditional error probabilities are small shows that their corresponding sample size $M$ must be large, so large in fact that it is asymptotically equivalent to its maximum value $N$.
@@ 2.7.3 Composite Hypotheses
For testing separated hypotheses $\theta\leqslant\theta_{0}$ vs. $\theta\geqslant\theta_{1}>\theta_{0}$ about the 1-dimensional parameter $\theta$ of an exponential family, Lorden (1983, Section 3) constructs an asymptotically optimal 3-stage test utilizing a description of the optimal stopping boundary related to Schwarz’s (1962) study of Bayes asymptotic shapes for fully sequential tests, described in Section 2.6. Let $n(\theta)$ denote the expected sample size to Schwarz’s boundary under $\theta$. Lorden’s test utilizes the “worst case” competing parameter value $\theta^{*}\in(\theta_{0},\theta_{1})$, which is characterized as the unique solution of (41) and which maximizes the expected sample size $n(\theta^{*})=\max_{\theta}n(\theta)\equiv n^{\star}$, given by (41). The first stage size of Lorden’s procedure is a fixed fraction of $n^{\star}$. If the procedure does not stop after the first stage, utilizing Schwarz’s boundary, the second stage brings the total sample size to $\min\{n^{\star},(1+\varepsilon)n(\widehat{\theta})\}$, where $\widehat{\theta}$ is the MLE of $\theta$ from the first stage data and $\varepsilon\searrow 0$ is a chosen sequence. Finally, if needed, the third stage brings the total sample size up to $n^{\star}$. Under (63), Lorden (1983, Theorem 1) proves that this test asymptotically minimizes the expected sample size to first order, not just for $\theta$ in the hypotheses but uniformly in $\theta$ over any interval in the parameter space containing $[\theta_{0},\theta_{1}]$. The first order term is of order $\log\alpha_{i}^{-1}$, as above, and the second order term is of order $O(((\log\alpha_{i}^{-1})\log\log\alpha_{i}^{-1})^{1/2})$, $i=0,1$.
@@ 3 Sequential Changepoint Detection: Lorden’s Minimax Change Detection Theory
In a variety of practical applications, the observed process abruptly changes statistical properties at an unknown point in time. Examples include but are not limited to aerospace navigation and flight systems integrity monitoring, cyber-security, identification of terrorist activity, industrial monitoring, air pollution monitoring, radar, sonar, and electrooptics surveillance systems. As a result, it has been of interest to many practitioners for some time. In classical quickest changepoint detection, one’s aim is to detect changes in the distribution as quickly as possible, minimizing the expected delay to detection assuming the change is in effect.
@@
Historically, the subject of changepoint detection first began to emerge in the 1920–1930s motivated by quality control considerations. Shewhart’s charts were popular in the past Shewhart (1931). Optimal and nearly optimal sequential detection procedures were developed much later in the 1950s–1970s, after the emergence of Sequential Analysis Wald (1947). The ideas set in motion by Shewhart and Wald have formed a platform for vast research on sequential changepoint detection.
The desire to detect the change quickly causes one to be trigger-happy, which, on one hand, leads to an unacceptably high level of false alarm rate – terminating the process prematurely before a real change has occurred. On the other hand, attempting to avoid false alarms too strenuously causes a long delay between the true change point and the time it is detected. Thus, the essence of the problem is to attain a tradeoff between two contradicting performance measures – the loss associated with the delay in detecting a true change and that associated with raising a false alarm. A good sequential detection procedure is expected to minimize the average loss associated with the detection delay, subject to a constraint on the loss associated with false alarms, or vice versa.
Let $p_{\nu}({\mathbf{X}}^{n})=p(X_{1},\dots,X_{n}|\nu)$ denote the joint probability density of the sample ${\mathbf{X}}^{n}=(X_{1},\dots,X_{n})$ when the changepoint $\nu$ is fixed ($0\leqslant\nu<\infty$) and $p_{\infty}({\mathbf{X}}^{n})$ the joint density when $\nu=\infty$, i.e., when there is never a change. Let ${\mathsf{P}}_{\nu},{\mathsf{P}}_{\infty}$ and ${\mathsf{E}}_{\nu},{\mathsf{E}}_{\infty}$ denote the corresponding probability measures and expectations. Assume that the observations $\{X_{n}\}_{n\geqslant 1}$ are independent and such that $X_{1},\ldots,X_{\nu}$ are each distributed according to a common (pre-change) density $f_{0}(x)$, while $X_{\nu+1},X_{\nu+2},\ldots$ each follows a common (post-change) density $f_{1}(x)$. In this i.i.d. case, the model can be written as
@@
Denote by $\operatorname{\mathcal{H}}_{\infty}:\nu=\infty$ the hypothesis that the change never occurs and by $\operatorname{\mathcal{H}}_{\nu}$ the hypothesis that the change occurs at time $0\leqslant\nu<\infty$. Define the likelihood ratio and the corresponding log-likelihood ratio for the $t$-th observation $X_{t}$
$\mathsf{LR}_{t}=\frac{f_{1}(X_{t})}{f_{0}(X_{t})},\quad Z_{t}=\log\frac{f_{1}(X_{t})}{f_{0}(X_{t})},\quad t=1,2,\dots$
We now introduce the CUMULATIVE SUM (CUSUM) detection procedure, which was first proposed by Page (1954). The changepoint detection problem can be considered as a problem of testing two hypotheses: $\operatorname{\mathcal{H}}_{\nu}$ that the change occurs at a fixed point $0\leqslant\nu<\infty$ against the alternative $\operatorname{\mathcal{H}}_{\infty}:\nu=\infty$ that the change never occurs. The likelihood ratio between these hypotheses is $\Lambda_{n}^{\nu}=\prod_{t=\nu+1}^{n}\mathsf{LR}_{t}$ for $\nu<n$ and $1$ for $\nu\geqslant n$. Since the hypothesis $\operatorname{\mathcal{H}}_{\nu}$ is composite, we may apply the GLR approach maximizing the likelihood ratio $\Lambda_{n}^{\nu}$ over $\nu$ to obtain the GLR statistic
$V_{n}=\max_{\nu\geqslant 0}\prod_{t=\nu+1}^{n}\mathsf{LR}_{t},\quad n\geqslant 1.$
It is easy to verify that the logarithmic version of this statistic
$W_{n}=\log V_{n}=\max_{\nu\geqslant 0}\sum_{t=\nu+1}^{n}Z_{t}$ (65)
obeys the recursion
@@
Page (1954) suggested measuring the risk due to a false alarm by the mean time to false alarm ${\mathsf{E}}_{\infty}[T]$ and the risk associated with a true change detection by the mean time to detection ${\mathsf{E}}_{0}[T]$ when the change occurs at the very beginning. He called these performance characteristics the Average Run Length (ARL). Page also analyzed the CUSUM procedure defined by (65)–(67) using these operating characteristics. While the false alarm rate is reasonable to measure by the ARL to false alarm $\operatorname{\mathsf{ARLFA}}(T)={\mathsf{E}}_{\infty}[T]$, the risk due to a true change detection is reasonable to measure by the conditional expected delay to detection ${\mathsf{E}}_{\nu}[T-\nu|T>\nu]$ for any possible change point $\nu\in\mathbb{Z}_{+}$ but not necessarily by the ARL to detection ${\mathsf{E}}_{0}[T]$. Ideally, a good detection procedure has to guarantee small values of the expected detection delay for all change points $\nu\in\mathbb{Z}_{+}$ when $\operatorname{\mathsf{ARLFA}}(T)$ is set at a certain level. However, if the false alarm risk is measured in terms of the ARL to false alarm, i.e., it is required that $\operatorname{\mathsf{ARLFA}}(T)\geqslant\gamma$ for some $\gamma\geqslant 1$, then a procedure that minimizes the conditional expected delay to detection ${\mathsf{E}}_{\nu}[T-\nu|T>\nu]$ uniformly over all $\nu$ does not exist. For this reason, we have to resort to different optimality criteria, e.g., Bayesian and minimax criteria.
@@
Lorden (1971) was the first who addressed the minimax change detection problem and developed the first minimax theory. He proposed to measure the false alarm risk by the ARL to false alarm $\operatorname{\mathsf{ARLFA}}(T)={\mathsf{E}}_{\infty}[T]$, i.e., to consider the class of change detection procedures
${\mathbb{C}}(\gamma)=\left\{T:\operatorname{\mathsf{ARLFA}}(T)\geqslant\gamma\right\},\quad\gamma\geqslant 1$
and the risk associated with detection delay by the worst-case expected detection delay
@@
In other words, the conditional expected detection delay is maximized over all possible trajectories of observations $(X_{1},\dots,X_{\nu})$ up to the changepoint and then over the changepoint $\nu$.
@@
Lorden (1971) showed that Page’s CUSUM procedure is first-order asymptotically minimax as $\gamma\to\infty$. This result was the first optimality result in the minimax change detection problem. Due to the importance of this result as well as the popularity of Lorden’s minimax criterion not only among statisticians but also among a variety of practitioners, we now present more details.
To establish the asymptotic optimality of Page’s CUSUM procedure, Lorden uses an interesting trick that allows one to use one-sided hypothesis tests for evaluation of a set of change detection procedures, including Page’s. Let $\tau=\tau(\alpha)$ be a stopping time with respect to $X_{1},X_{2},\ldots$ such that
@@
The following theorem (see Theorem 2 in Lorden (1971)) allows us to construct nearly optimal change detection procedures and to prove near optimality of the CUSUM procedure. Recall that ${\mathsf{P}}_{\infty}$ corresponds to the distribution with density $f_{0}(x)$ and ${\mathsf{P}}_{0}$ corresponds to the distribution with density $f_{1}(x)$.
@@
The cumulative LLR for the sample $(X_{k+1},\dots,X_{n})$ is $\lambda_{n}^{k}=\sum_{t=k+1}^{n}Z_{t}.$ Let $\tau(\alpha)$ denote the stopping time of the one-sided SPRT for testing $f_{0}$ versus $f_{1}$ with threshold $|\log\alpha|$:
$\tau(\alpha)=\inf\left\{n\geqslant 1:\lambda_{n}^{0}\geqslant|\log\alpha|\right\}.$
Then ${\mathsf{P}}_{\infty}(\tau(\alpha)<\infty)\leqslant\alpha$, so condition (70) holds. If the Kullback-Leibler information number $I={\mathsf{E}}_{0}[Z_{1}]$ is positive and finite, then it is well-known that
@@
It follows from Theorem 7 that if we set $\alpha=\gamma^{-1}$, then ${\mathsf{E}}_{\infty}[T_{a_{\gamma}}]\geqslant\gamma$, so $T_{a_{\gamma}}\in{\mathbb{C}}(\gamma)$, and
@@
In Theorem 3 Lorden (1971) proved this fact using a quite sophisticated argument. Note, however, that Lai (1998) established the lower bound (73) in a general non-i.i.d. case assuming that $n^{-1}\lambda_{\nu+n}^{\nu}$ converges as $n\to\infty$ to a positive and finite number $I$ under a certain additional condition. In the i.i.d. case, by the SLLN $n^{-1}\lambda_{\nu+n}^{\nu}$ converges to the Kullback–Leibler information number $I$ almost surely under ${\mathsf{P}}_{\nu}$, which implies that as $M\to\infty$ for all $\varepsilon>0$
@@
Using (74), the lower bound (73) can be obtained from Theorem 1 in Lai (1998) as a particular case.
To handle a composite parametric post-change hypothesis, which is typically the case in applications, let $f_{\theta}(x)$ be the post-change density, $\theta\in\Theta$. Write $Z_{n}(\theta)=\log[f_{\theta}(X_{n})/f_{0}(X_{n})]$. Then inequality (72) in Theorem 7 holds for expectation ${\mathsf{E}}_{\theta}[\tau^{*}]$. Also, if we assume that the Kullback-Leibler information number $I(\theta)={\mathsf{E}}_{\theta}[Z_{1}(\theta)]$ is positive and finite, then asymptotic lower bound (73) holds with $I(\theta)$, i.e.,

### v2

@@ 2.4 Near Uniform Optimality of the GLR SPRT for Composite Hypotheses
and where $\zeta(\theta,\theta_{i})$, $i=0,1$, are defined in (39)–(40). These approximations hold significance for frequentist problems, which are typically of primary interest in most applications. While there are no strict upper bounds on the error probabilities, leading to no specific prescription on how to embed the GSLRT into class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$, the asymptotic approximations (44) enable us to select thresholds $a_{i}$ in the stopping times $\widehat{T}_{i}$ so that $\hat{\alpha}_{i}(\theta_{i})\approx\alpha_{i}$, $i=0,1$, at least for sufficiently small $\alpha_{i}$. Note that in this latter case, the threshold $a$ in (41) should be replaced with $a_{i}$, the roots of the transcendental equations
@@
where the $o(1)$ term is of order $O(\log|\log\alpha_{\max}|/|\log\alpha_{\max}|)$. It’s noteworthy that this result holds true not only in the asymptotically symmetric case where $\log\alpha_{0}\sim\log\alpha_{1}$ and $a_{0}\sim a_{1}$ as $\alpha_{\max}\to 0$, but also in the asymmetric case where $a_{0}$ and $a_{1}$ diverge with different rates, as long as $a_{1}e^{-a_{0}}\to 0$.
Note that the Schwarz–Lorden asymptotic theory operates under the assumption of a fixed indifference zone that does not permit local alternatives, meaning that $\theta_{1}$ cannot approach $\theta_{0}$ as $c\to 0$. In simpler terms, this theory is confined to scenarios where the width of the indifference zone $\theta_{1}-\theta_{0}$ is considerably larger than $c^{1/2}$.
We conclude this section by mentioning Lorden’s (1973) paper on the properties of the one-sided (open-ended) GSLRTs for the one-parameter exponential family. These tests reject a null hypothesis $\theta=\theta_{0}$ in favor of $\theta>\theta_{0}$ within the class of stopping times satisfying ${\mathsf{P}}_{\theta_{0}}(T<\infty)\leqslant\alpha$ for a prescribed $30<\alpha<1/$ 3.
@@ 2.5 Optimal Multistage Testing
Beginning with the simple vs. simple testing setup of Section 2.1.1 and adopting the notation there, some of Lorden’s main ideas can be seen by first considering the symmetric case where the error probabilities $\alpha_{0},\alpha_{1}\to 0$ in such a way that
@@
$\frac{1/n}{(1-t/n)^{2}}=\frac{n}{\delta_{t}^{2}}=\frac{t+\delta_{t}}{\delta_{t}^{2}}=o(1)\quad\mbox{as $t\to\infty$.}$
A similar argument shows that the other probability in (46) approaches 00 as well.
In this symmetric situation, an optimal 2-stage competitor to the SPRT can be described in terms of $n(t)$, which is the size of the first stage with $t$ taken to be the larger of the two sides of (45). Note that, for either $i=0$ or $1$, we have $tI_{i}\sim\log\alpha_{1-i}^{-1}$ so that $t$ is asymptotically the same as the expected stopping time of the SPRT (under either hypothesis) and the first stage $n(t)$ is of the same order but slightly larger. The procedure stops after the first stage if
@@
Lorden shows that no 2-stage test can be asymptotically optimal, itself a nontrivial result that we discuss in the next section. For this case Lorden gives a 3-stage procedure that is a slight modification of the one above. Letting $t_{1}$ and $t_{2}$ be the left- and right-hand sides of (45), respectively, the first stage of the procedure is of size $\min\{n(t_{1}),n(t_{2})\}$, and the second stage (if needed) brings the total sample size to $\max\{n(t_{1}),n(t_{2})\}$, both using the stopping rule (48) and corresponding decision rule. If not stopped by the second stage, a third stage brings the total sample size to that of the fixed-sample size with error probabilities $\alpha_{i}$, which is $\leqslant C\max\{n(t_{1}),n(t_{2})\}$ as above, and uses that terminal decision rule. Since $n(t_{i+1})\sim t_{i+1}\sim\log\alpha_{1-i}^{-1}/I_{i}$ for both $i=0,1$, and (49) and (50) hold for $n(t_{1})$ and $n(t_{2})$, respectively, the expected sample size of this 3-stage procedure is asymptotically equal to the corresponding side of (45), and is thus minimized under both the null and alternative.
@@ 2.5.2 The Necessity of 3 Stages
For testing $f_{0}$ vs. $f_{1}$ in the setup of Section 2.1.1, let $N$ denote the sample size of a $k$-stage ($k\geqslant 2$) test with error probabilities $\alpha_{0}$ and $\alpha_{1}$, and let $M$ be the total sample size of this test after $k-1$ stages. If $N$ is asymptotically optimal as $\alpha_{0},\alpha_{1}\to 0$ and $g$ is a density distinct from $f_{0}$ such that
@@
$M\to\frac{\log\alpha_{0}^{-1}}{I(g,f_{0})}\quad\mbox{in $g$-probability, and}\quad{\mathsf{E}}_{g}[M]\sim\frac{\log\alpha_{0}^{-1}}{I(g,f_{0})}\sim{\mathsf{E}}_{g}[N]\quad\mbox{as $\alpha_{0}\to 0$.}$
Lorden’s proof of this theorem is technical and requires detailed upper bounds on the conditional error probabilities after the $(k-1)$st stage; that is, the probabilities of test error given the first $M$ observations. Roughly speaking, showing that these error probabilities are small shows that their corresponding sample size $M$ must be large, so large in fact that it is asymptotically equivalent to its maximum value $N$.
@@ 2.5.3 Composite Hypotheses
For testing separated hypotheses $\theta\leqslant\theta_{0}$ vs. $\theta\geqslant\theta_{1}>\theta_{0}$ about the 1-dimensional parameter $\theta$ of an exponential family, Lorden (1983, Section 3) constructs an asymptotically optimal 3-stage test utilizing a description of the optimal stopping boundary related to Schwarz’s (1962) study of Bayes asymptotic shapes for fully sequential tests, described in Section 2.4. Let $n(\theta)$ denote the expected sample size to Schwarz’s boundary under $\theta$. Lorden’s test utilizes the “worst case” competing parameter value $\theta^{*}\in(\theta_{0},\theta_{1})$ which maximizes the expected sample size $n(\theta^{*})=\max_{\theta}n(\theta)\equiv n^{\star}$. The first stage size of Lorden’s procedure is a fixed fraction of $n^{\star}$. If the procedure does not stop after the first stage, utilizing Schwarz’s boundary, the second stage brings the total sample size to $\min\{n^{\star},(1+\varepsilon)n(\widehat{\theta})\}$, where $\widehat{\theta}$ is the MLE of $\theta$ from the first stage data and $\varepsilon\searrow 0$ is a chosen sequence. Finally, if needed, the third stage brings the total sample size up to $n^{\star}$. Under (51), Lorden (1983, Theorem 1) proves that this test asymptotically minimizes the expected sample size to first order, not just for $\theta$ in the hypotheses but uniformly in $\theta$ over any interval in the parameter space containing $[\theta_{0},\theta_{1}]$. The first order term is of order $\log\alpha_{i}^{-1}$, as above, and the second order term is of order $O(((\log\alpha_{i}^{-1})\log\log\alpha_{i}^{-1})^{1/2})$, $i=0,1$.
@@ 3 Sequential Changepoint Detection: Lorden’s Minimax Change Detection Theory
In numerous practical applications, the observed process undergoes an abrupt change in statistical properties at an unknown point in time. Examples encompass aerospace navigation and flight systems integrity monitoring, cyber-security, identification of terrorist activity, industrial monitoring, air pollution monitoring, radar, sonar, and electrooptics surveillance systems. Consequently, this problem has garnered interest from many practitioners for some time.
In classical quickest changepoint detection, the objective is to detect changes in the distribution as swiftly as possible, thereby minimizing the expected delay to detection assuming the change is in effect.
@@
Historically, the field of changepoint detection began to take shape in the 1920s to 1930s, spurred by considerations in quality control. Shewhart’s charts were particularly influential during this period (Shewhart 1931). However, optimal and nearly optimal sequential detection procedures didn’t come into prominence until much later, in the 1950s to 1970s, following the advent of Sequential Analysis (Wald 1947). The concepts initiated by Shewhart and Wald laid the foundation for extensive research into sequential changepoint detection.
The desire to detect the change quickly often leads to being “trigger-happy,” which, on one hand, results in an unacceptably high false alarm rate – terminating the process prematurely before a real change has occurred. On the other hand, attempting to avoid false alarms too strenuously causes a long delay between the true change point and its detection. Thus, the essence of the problem lies in achieving a tradeoff between two conflicting performance measures – the loss associated with the delay in detecting a true change and that associated with raising a false alarm. An efficient detection procedure is expected to minimize the average loss associated with the detection delay, while subject to a constraint on the loss associated with false alarms, or vice versa.
Let $p_{\nu}({\mathbf{X}}^{n})=p(X_{1},\dots,X_{n}|\nu)$ denote the joint probability density of the sample ${\mathbf{X}}^{n}=(X_{1},\dots,X_{n})$ when the changepoint $\nu$ is fixed ($0\leqslant\nu<\infty$) and $p_{\infty}({\mathbf{X}}^{n})$ the joint density when $\nu=\infty$, i.e., when there is never a change. Let ${\mathsf{P}}_{\nu},{\mathsf{P}}_{\infty}$ and ${\mathsf{E}}_{\nu},{\mathsf{E}}_{\infty}$ denote the corresponding probability measures and expectations. Assume that the observations $\{X_{n}\}_{n\geqslant 1}$ are independent and such that $X_{1},\ldots,X_{\nu}$ are each distributed according to a common (pre-change) density $f_{0}(x)$, while $X_{\nu+1},X_{\nu+2},\ldots$ each follows a common (post-change) density $f_{1}(x)$. Hence, the model can be represented as
@@
Denote by $\operatorname{\mathcal{H}}_{\infty}:\nu=\infty$ the hypothesis that the change never occurs and by $\operatorname{\mathcal{H}}_{\nu}$ the hypothesis that the change occurs at time $0\leqslant\nu<\infty$. Let $Z_{t}=\log[f_{1}(X_{t})/f_{0}(X_{t})]$ denote the LLR for the $t$-th observation $X_{t}$.
We now introduce the CUMULATIVE SUM (CUSUM) detection procedure, which was first proposed by Page (1954). The changepoint detection problem can be viewed as a problem of testing two hypotheses: $\operatorname{\mathcal{H}}_{\nu}$ that the change occurs at a fixed point $0\leqslant\nu<\infty$ against the alternative $\operatorname{\mathcal{H}}_{\infty}:\nu=\infty$ that the change never occurs. The LLR between these hypotheses is $\lambda_{n}^{\nu}=\sum_{t=\nu+1}^{n}Z_{t}$ for $\nu<n$ and 00 for $\nu\geqslant n$. Since the hypothesis $\operatorname{\mathcal{H}}_{\nu}$ is composite, we may employ the GLR approach, maximizing the LLR $\lambda_{n}^{\nu}$ over $\nu$, to obtain the log-GLR statistic:
$W_{n}=\max_{\nu\geqslant 0}\sum_{t=\nu+1}^{n}Z_{t},$ (53)
which follows the recursion
@@
Page (1954) proposed measuring the risk due to a false alarm by the mean time to false alarm ${\mathsf{E}}_{\infty}[T]$ and the risk associated with a true change detection by the mean time to detection ${\mathsf{E}}_{0}[T]$ when the change occurs at the very beginning. These are commonly known as the Average Run Length (ARL). Page also analyzed the CUSUM procedure defined by equations (53)–(55) using these operating characteristics.
While the false alarm rate is reasonable to measure by the ARL to false alarm $\operatorname{\mathsf{ARLFA}}(T)={\mathsf{E}}_{\infty}[T]$, the risk due to a true change detection is better measured by the conditional expected delay to detection ${\mathsf{E}}_{\nu}[T-\nu|T>\nu]$ for any possible change point $\nu\in\mathbb{Z}_{+}$, rather than by the ARL to detection ${\mathsf{E}}_{0}[T]$. Ideally, a good detection procedure should guarantee small values of the expected detection delay for all change points $\nu\in\mathbb{Z}_{+}$ when $\operatorname{\mathsf{ARLFA}}(T)$ is set at a certain level. However, if the false alarm risk is measured in terms of the ARL to false alarm, i.e., it is required that $\operatorname{\mathsf{ARLFA}}(T)\geqslant\gamma$ for some $\gamma\geqslant 1$, then a procedure that minimizes the conditional expected delay to detection ${\mathsf{E}}_{\nu}[T-\nu|T>\nu]$ uniformly over all $\nu$ does not exist. For this reason, we must resort to different optimality criteria, such as Bayesian and minimax criteria.
@@
Lorden (1971) was the first who addressed the minimax change detection problem and developed the first minimax theory. He proposed to measure the false alarm risk by the ARL to false alarm $\operatorname{\mathsf{ARLFA}}(T)={\mathsf{E}}_{\infty}[T]$, i.e., to consider the class of change detection procedures ${\mathbb{C}}(\gamma)=\left\{T:\operatorname{\mathsf{ARLFA}}(T)\geqslant\gamma\right\}$ for some $\gamma\geqslant 1$, and the risk associated with detection delay by the worst-case expected detection delay
@@
In other words, the conditional expected detection delay is maximized over all possible trajectories $(X_{1},\dots,X_{\nu})$ up to the changepoint and then over the changepoint $\nu$.
@@
Lorden (1971) demonstrated that Page’s CUSUM procedure achieves first-order asymptotic minimax optimality as $\gamma$ approaches infinity. This groundbreaking finding marked the initial optimality result in the minimax change detection problem. Given the significance of this outcome and the widespread adoption of Lorden’s minimax criterion not only within statistical circles but also across various practical domains, we proceed to provide further elaboration.
To establish the asymptotic optimality of Page’s CUSUM procedure, Lorden employs an intriguing method that permits the utilization of one-sided hypothesis tests to assess a collection of change detection procedures, among them Page’s method. Let $\tau=\tau(\alpha)$ be a stopping time with respect to $X_{1},X_{2},\ldots$ such that
@@
The subsequent theorem, resembling Theorem 2 in Lorden (1971), empowers the construction of nearly optimal change detection procedures and facilitates the demonstration of the near optimality of the CUSUM procedure. It’s important to recall that ${\mathsf{P}}_{\infty}$ denotes the distribution characterized by the density $f_{0}(x)$, while ${\mathsf{P}}_{0}$ corresponds to the distribution with density $f_{1}(x)$.
@@
The cumulative LLR for the sample $(X_{k+1},\dots,X_{n})$ is $\lambda_{n}^{k}=\sum_{t=k+1}^{n}Z_{t}.$ Let $\tau(\alpha)=\inf\left\{n\geqslant 1:\lambda_{n}^{0}\geqslant|\log\alpha|\right\}$ denote the stopping time of the one-sided SPRT for testing $f_{0}$ versus $f_{1}$ with threshold $|\log\alpha|$. Then ${\mathsf{P}}_{\infty}(\tau(\alpha)<\infty)\leqslant\alpha$, so condition (58) holds. If the Kullback-Leibler information number $I={\mathsf{E}}_{0}[Z_{1}]$ is positive and finite, then it is well-known that
@@
It follows from Theorem 6 that setting $\alpha=\gamma^{-1}$ gives ${\mathsf{E}}_{\infty}[T_{a_{\gamma}}]\geqslant\gamma$, so $T_{a_{\gamma}}\in{\mathbb{C}}(\gamma)$, and
@@
Theorem 3 of Lorden (1971) establishes this fact using a rather sophisticated argument. Note, however, that Lai (1998) established the lower bound (61) in a general non-i.i.d. case, assuming that $n^{-1}\lambda_{\nu+n}^{\nu}$ converges to a positive and finite number $I$ as $n\to\infty$, under a certain additional condition. In the i.i.d. case, by the SLLN $n^{-1}\lambda_{\nu+n}^{\nu}$ converges to the Kullback–Leibler information number $I$ almost surely under ${\mathsf{P}}_{\nu}$. This implies that as $M\to\infty$ for all $\varepsilon>0$
@@
Using (62), the lower bound (61) can be obtained from Theorem 1 in Lai (1998).
To handle a composite parametric post-change hypothesis, which is typical in many applications, let $f_{\theta}(x)$ be the post-change density, where $\theta\in\Theta$. Denote $Z_{n}(\theta)=\log[f_{\theta}(X_{n})/f_{0}(X_{n})]$. Then, inequality (60) in Theorem 6 holds for expectation ${\mathsf{E}}_{\theta}[\tau^{*}]$. Additionally, assuming that the Kullback-Leibler information number $I(\theta)={\mathsf{E}}_{\theta}[Z_{1}(\theta)]$ is positive and finite, then asymptotic lower bound (61) holds with $I(\theta)$, i.e.,